OpenStudy (anonymous):

graph each system of inequalities 2x-y <5 3x + y > 4 -

OpenStudy (phi):

this is more algebra. first put them into the form y> or < mx+b (btw use >= if you can't type ≥)

OpenStudy (phi):

It's a bit trickier than = because if you multiply both sides by -1 the inequality switches.

OpenStudy (anonymous):

ok well can you help me

OpenStudy (phi):

2x-y <5 add y to both sides

OpenStudy (anonymous):

y is nothing

OpenStudy (phi):

You don't like letters. y is y. You can add y's or subtract y's. That is what algebra is all about. So go with the flow here.

OpenStudy (anonymous):

ok lol.. help me please im exhausted

OpenStudy (phi):

add just means write down +y for example: add y to 3: 3+y

OpenStudy (phi):

2x-y <5 add y to both sides

OpenStudy (anonymous):

3y

OpenStudy (phi):

do this (pay attention) 2x-y+y < 5+y (I added y to both sides)

OpenStudy (phi):

now -y+y= 0 (no matter what y is when we subtract them it is zero)

OpenStudy (anonymous):

ok

OpenStudy (phi):

so simplify 2x-y+y to just 2x. We now have 2x<5+y add -5 to both sides.

OpenStudy (anonymous):

addd -5 to 2?

OpenStudy (phi):

to 2x (not just 2) remember if the term is not a simple number, you just write down -5+2x

OpenStudy (anonymous):

so -3x

OpenStudy (phi):

you can't combine numbers with terms like 3x. So you write down -5+2x or 2x-5

OpenStudy (anonymous):

ok then what

OpenStudy (phi):

what do we have now?

OpenStudy (anonymous):

idk im so exhausted

OpenStudy (anonymous):

wait so i graph that ?

OpenStudy (phi):

we have -5+2x < y This says y is bigger than 2x-5 now you know we can re-write this as y > 2x - 5 This says y is bigger than 2x -5 Now we plot the line y= 2x -5 . We make it dotted because the inequality is > (not ≥) then we shade the area above the line.

OpenStudy (phi):

y= 2x-5 is a line y> 2x-5 is all those places where y is above (bigger than the y that is on the line)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so go down to -5

OpenStudy (anonymous):

up 2 over 1

OpenStudy (phi):

yes.

OpenStudy (anonymous):

ok thats one line

OpenStudy (anonymous):

can you help me with this one 3x+y > 9 2x + y >= 6

OpenStudy (phi):